What is the probability of making exactly 3 of 7 free throws when a player shoots 53%.
Is it N!/(N-r)! or 7! / 4! * .53 = or 80 % chance of making excacly 3 of 7 free throws.

Answer by stanbon(75887)

What is the probability of making exactly 3 of 7 free throws when a player shoots 53%.
----
Binomial Problem with n = 7 ; p(score) = 0.53
----
P(x = 3) = 7C3*0.53^3*0.47^4 = 35*0.149*0.0488 = 0.2543
========================
Cheers,
Stan H.